a 10L helium balloon at a pressure of 2 bar at 27C. some O2 is now in balloon and final volume is made 30L. the pressure also fell to 1 bar at same temperature. the partial pressure of O2 at end is?

Grade:12

3 Answers

Akash Kumar Dutta
11 years ago

Dear Himanshu,

for helium.
n(He)=P(1)V(1)/RT.
for oxygen n(O)=P(2)V(2)/RT

partial pressure of O2= (mole fraction of O2).(Total pressure)
= [  P(2).V(2) / P(2).V(2) + P(1).V(1)  ].(1)
= 1.30 / 1.30 + 2.10
=30/30+20=3/5 bar
Hence 0.6 bar (ANS)

Regards.

lokesh soni
11 years ago

p1v1=p2v2

let partial pressure of helium in 30L be P

then p1*30=2*10

p1=20/30 bar=2/3bar

partial pressure of O2=p2

p1+p2=1bar

partial pressure of O2=1/3 bar

Akash Kumar Dutta
11 years ago

Dear Himanshu,

I am really sorry!
ACTUALLY I MADE A MISTAKE IN THE ANS.
SO HERES ANOTHER ATTEMPT:-

For Helium,
P(1)=2 bar,V(1)=10 L
P(2)= x (let)  , V (2)= 30 L
Hence, P(2)=x=2/3 bar.      [ BY APPLYING BOYLES LAW. ]

So Pressure of oxygen=1-2/3=1/3 bar = .33 bar (ANS).
HOPE THIS HELPS.

Regards.